OpenStudy (anonymous):

find the quotient (-25c^5-5c^2) / 5c

OpenStudy (anonymous):

\[\LARGE \frac{-25c^5-5c^2}{5c}=\frac{5c(-5c^4-c)}{5c}=-5c^4-c\] I think you should know it now... :)